An ideal gas undergoes a reversible isothermal exp
Answer #1

Isothermal reversible expansion condition at T = 57 °C = 330 K,

Work done, W = -nRT ln(V2 / V1)

Entropy change,   dS = dQ / T

We know that,   dU = dQ - dW

   => dQ = dW = -nRT ln(V2 / V1) ( since dU = 0 for isothermal condition )

Clubbing above equation with the entropy change equation,

dS = dQ / T = nRT ln(V2 / V1) / T

Putting the appropriate values from the question,

dS = nRT ln(V2 / V1) / T

=> 36 = n*8.314*330 * ln(4.5 / 1.5) / 330

=>   n = 3.941 moles ~ 4 moles